# Bayesian Doubly Robust Causal Inference via Posterior Coupling [^1] October 16, 2025

### v2

@@ Results
To estimate the ATE (i.e., the risk difference), we applied Algorithm 4 described in Section 7.2. In this analysis, we define the subset of selected covariates as $S=\{j:|\bar{\bm{\alpha}}_{j}|\geq 0.01\}$, where $\bar{\bm{\alpha}}_{j}$ denotes the posterior mean in Step 1 of Algorithm 4 for the $j$-th covariate. From Step 1, 44 of the 58 covariates are selected. Note that there are 58 candidates because dummy variables are included. The posterior mean of the ATE and its 95% CI for the G-formula is $0.229$ $(0.032,0.395)$, with a CI length of $0.363$. Without applying the algorithm (i.e., without covariate selection), the posterior mean of the ATE and its 95% CI is $0.043$ (0.075,0.205)(-0.075,0.205), with a CI length of $0.280$. In contrast, with the algorithm, the estimates are $0.188$ $(0.122,0.312)$, with a CI length of $0.190$. These results suggest that applying the algorithm allows more accurate estimation of causal effects, which is consistent with the discussion in Appendix D. Based on these findings, we recommend that updating the selected posterior with ET is necessary to accurately estimate causal effects when there are many potential confounders.
@@ Discussion
In this manuscript, we propose a novel Bayesian doubly robust estimator whose posterior distribution can be described explicitly. Our proposed method achieves this by using an entropic tilting condition, which is related to the doubly robust estimator proposed by Bang and Robins (2005). This condition plays a role in modifying the posterior distribution for the outcome model by incorporating information from the propensity score model. As shown in both the mathematical discussions and the simulation results, our proposed method exhibits double robustness. The explicit posterior formulation also enables important extensions such as sensitivity analysis for unmeasured confounding and confounding selection in high-dimensional settings.
As mentioned in the Introduction, many Bayesian doubly robust estimators have been proposed (Saarela et al., 2016; Yiu et al., 2020; Antonelli et al., 2022; Breunig et al., 2025). However, an explicit description of the posterior distribution (i.e., prior distributions and likelihoods) is particularly attractive. The explicit posterior formulation enables various flexible estimation strategies, including hierarchical Bayesian modeling for clustered data—a structure appealing in several fields, including biostatistics and medicine (Yang, 2018; Chang and Stuart, 2022). Additionally, combining dynamic treatment regime settings with Bayesian doubly robust inference is attractive (Dong et al., 2025). Furthermore, as mentioned in Appendix E, the posterior description enables the construction of an algorithm for estimating the optimal number of strata for propensity score subclassification (Orihara and Momozaki, 2024).
Acknowledgement: We have prepared an R package for our proposed methodology. For more details, see the GitHub page of Dr. Momozaki: https://github.com/t-momozaki/DRBayes. We would like to express our gratitude to Dr. Yu Luo for their comments and supports.
@@ Appendix B Proofs
First, considering the first two components of (B.4) (for $a=1$):
@@
Here, the last equation is a Taylor expansion with respect to $\bm{\alpha}^{0}$ around $\bm{\alpha}$. From the same discussion, the last two components of (B.4) (for $a=0$) becomes
@@ Appendix C Benefits of incorporating entropic tilting
Thus, the proposed Bayesian DR method is clearly advantageous in terms of convergence rate reduction if the propensity score model is correctly specified. Additionally, from Lemma 1, it is expected that the posterior distribution for the outcome model becomes the same both with and without using ET if the propensity score model is misspecified. These points are also confirmed by the simulation experiments presented in the main manuscript.
@@ Appendix D Additional simulation experiments
Appendix E Entropic tilting using propensity score subclassification
@@
$\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{K}\left(\frac{A_{i}}{n_{k1}/n_{k+}}-\frac{(1-A_{i})}{1-n_{k1}/n_{k+}}\right)Y_{i}{\rm I}_{\{\hat{c}_{k-1}\leq\hat{e}_{i}<\hat{c}_{k}\}},$ (E.1)
@@
Compared with (2.2), $n_{k1}/n_{+}$ in (E.1) can be viewed as the propensity score for each stratum. Therefore, modifying the ET condition (3.4):
@@
$B_{n}^{Sub}(\bm{\alpha},\bm{\beta})=\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{K}\left(\frac{A_{i}}{n_{k1}/n_{+}}-\frac{1-A_{i}}{1-n_{k1}/n_{+}}\right)\left(Y_{i}-m_{A_{i}}(\bm{X}_{i};\bm{\beta})\right){\rm I}_{k}(\bm{X}_{i}).$ (E.2)
In fact, using (E.2), the proposed Bayesian procedure also achieves (approximately) double robustness, based on the same discussion presented in the main manuscript.

### v3

@@ Results
To estimate the ATE (i.e., the risk difference), we applied Algorithm 4 described in Section 7.2. In this analysis, we define the subset of selected covariates as $S=\{j:|\bar{\bm{\alpha}}_{j}|\geq 0.01\}$, where $\bar{\bm{\alpha}}_{j}$ denotes the posterior mean in Step 1 of Algorithm 4 for the $j$-th covariate. From Step 1, 44 of the 58 covariates are selected. Note that there are 58 candidates because dummy variables are included. The posterior mean of the ATE and its 95% CI for the G-formula is $0.229$ $(0.032,0.395)$, with a CI length of $0.363$. Without applying the algorithm (i.e., without covariate selection), the posterior mean of the ATE and its 95% CI is $0.043$ (0.075,0.205)(-0.075,0.205), with a CI length of $0.280$. In contrast, with the algorithm, the estimates are $0.188$ $(0.122,0.312)$, with a CI length of $0.190$. These results suggest that applying the algorithm allows more accurate estimation of causal effects, which is consistent with the discussion in Appendix D.1. Based on these findings, we recommend that updating the selected posterior with ET is necessary to accurately estimate causal effects when there are many potential confounders.
@@ Discussion
In this manuscript, we propose a novel Bayesian DR estimator whose posterior distribution can be described explicitly. Our proposed method achieves this by using an entropic tilting condition, which is related to the DR estimator proposed by Bang and Robins (2005). This condition plays a role in modifying the posterior distribution for the outcome model by incorporating information from the propensity score model. As shown in both the mathematical discussions and the simulation results, our proposed method exhibits double robustness. The explicit posterior formulation also enables important extensions such as sensitivity analysis for unmeasured confounding and confounding selection in high-dimensional settings.
As mentioned in the Introduction, many Bayesian DR estimators have been proposed (Saarela et al., 2016; Yiu et al., 2020; Antonelli et al., 2022; Breunig et al., 2025). However, an explicit description of the posterior distribution (i.e., prior distributions and likelihoods) is particularly attractive. The explicit posterior formulation enables various flexible estimation strategies, including hierarchical Bayesian modeling for clustered data—a structure appealing in several fields, including biostatistics and medicine (Yang, 2018; Chang and Stuart, 2022). Additionally, combining dynamic treatment regime settings with Bayesian DR inference is attractive (Dong et al., 2025). Furthermore, as mentioned in Appendix F, the posterior description enables the construction of an algorithm for estimating the optimal number of strata for propensity score subclassification (Orihara and Momozaki, 2024).
Acknowledgement: We have prepared an R package for our proposed methodology. For more details, see the GitHub page of Dr. Momozaki: https://github.com/t-momozaki/DRBayes. We would like to express our gratitude to Dr. Kosuke Inoue, Dr. Kuan Liu, and Dr. Yu Luo for their comments and support. Dr. Liu also provided her idea for the discussion in Appendix E.
@@ Appendix B Proofs
First, considering the first two components of (B.4) (i.e., for $a=1$):
@@
Here, the last equation is a Taylor expansion with respect to $\bm{\alpha}^{0}$ around $\bm{\alpha}$. From the same discussion, the last two components of (B.4) (i.e., for $a=0$) becomes
@@ Appendix C Benefits of incorporating entropic tilting
Thus, it is expected that the proposed Bayesian DR method is clearly advantageous in terms of convergence rate reduction when the propensity score model is correctly specified. Additionally, from Lemma 1, it is expected that the posterior distribution for the outcome model becomes the same both with and without using ET asymptotically, even if the propensity score model is misspecified. These points are also confirmed by the simulation experiments presented in the main manuscript.
@@ Appendix D Additional simulation experiments
Kang and Schafer’s misspecification setting
In this section, we additionally consider the situation in which the outcome model is misspecified following the misspecification manner in Kang and Schafer (2007). Specifically, we use the following misspecified covariates for the outcome model:
$\left(\exp\left\{\frac{X_{1}}{2}\right\},10+\frac{X_{2}}{1+\exp\left\{X_{1}\right\}},\left(0.6+\frac{X_{1}X_{3}}{25}\right)^{3},\left(20+X_{1}+X_{4}\right)^{2}\right)$
Here, we use the covariates after standardization. The other settings are the same as those in the simulation experiments in the main manuscript. As is obvious, this misspecification is severe and extreme compared with that in the main manuscript (Robins et al., 2007).
The results are summarized in Table 3; Bang and Robins’s DR estimator and the G-formula show similar tendencies in the main manuscript. Our proposed methods decrease bias compared with the G-formula; however, the CP exhibits undercoverage. The pruning algorithm can further improve bias, RMSE, and CP; however, it still exhibits undercoverage. The main results and this additional simulation experiment show that our proposed methodology can improve bias, at least compared with the G-formula; however, CP depends on the misspecification pattern of the outcome model.
Table 3: Summary of causal effect estimates under outcome model misspecification: The number of iteration is $2000$ and the true ATE is 110. The absolute bias (ABias), empirical standard error (ESE), root mean squared error (RMSE), coverage probability (CP), and average length of credible interval (AvL) of the estimated causal effects across 2000 iterations are summarized by estimation method (“Method” column).
Method Sample size: $n=500$
ABias ESE RMSE CP AvL
DR 0.019 1.266 1.266 100 7.103
G-formula 3.439 1.494 3.750 37.7 5.927
Proposed 1.565 1.092 1.909 88.8 5.879
Proposed
0.327 1.333 1.372 91.3 4.640
DR: Ordinaly non-Bayesian DR estimator that is asymptotically equivalent to Bang and Robins (2005); G-formula: Bayesian G-formula based method discussed in Daniels et al. (2023); Pruning: Using sample pruning algorithm for our proposed method described in Section 5.3.1
For our proposed method, sequential Monte Carlo with replacement is conducted. For all Bayesian methods, 20,000 posterior draws are used.
Appendix E Importance weighting with a single sensitivity parameter
When $\xi_{i}\equiv\xi$, we can consider another option for the sampling algorithm for $\xi$. Let $f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta},\xi)$ be the corresponding negative log-likelihood function of the outcome model including $\xi$. Then, the joint posterior of $(\bm{\alpha},\bm{\beta})$ is proportional to
$\pi(\bm{\alpha})\pi(\bm{\beta})\prod_{i=1}^{n}\exp\left\{-f(A_{i}\mid\bm{X}_{i};\bm{\alpha})\right\}\int\exp\left\{-f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta},\xi)\right\}g(\xi)d\xi$
$=\pi(\bm{\alpha})\pi(\bm{\beta})\exp\left\{-\sum_{i=1}^{n}f(A_{i}\mid\bm{X}_{i};\bm{\alpha})\right\}\int\exp\left\{-\sum_{i=1}^{n}f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta},\xi)\right\}g(\xi)d\xi,$ (E.1)
where $g(\xi)$ is a density function of $\xi$. In the same manner as in the main manuscript, the posterior samples from the above posterior can be generated by an importance sampling with important weight proportional to
$\frac{\int\exp\left\{-\sum_{i=1}^{n}f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta},\xi)\right\}g(\xi)d\xi}{\exp\left\{-\sum_{i=1}^{n}f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta})\right\}}.$
Assuming that random sample generation from $g(\cdot)$ is tractable, we propose the following procedure to obtain the posterior sample from (E.1).
Algorithm A. 1.
$w^{(s)}\equiv\frac{M^{-1}\sum_{m=1}^{M}\exp\left\{-\sum_{i=1}^{n}f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta}^{(s)},\xi^{(m)})\right\}}{\exp\left\{-\sum_{i=1}^{n}f(Y_{i}\mid A_{i},\bm{X}_{i};\bm{\beta}^{(s)})\right\}}.$
Appendix F Entropic tilting using propensity score subclassification
@@
$\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{K}\left(\frac{A_{i}}{n_{k1}/n_{k+}}-\frac{(1-A_{i})}{1-n_{k1}/n_{k+}}\right)Y_{i}{\rm I}_{\{\hat{c}_{k-1}\leq\hat{e}_{i}<\hat{c}_{k}\}},$ (F.1)
@@
Compared with (2.2), $n_{k1}/n_{+}$ in (F.1) can be viewed as the propensity score for each stratum. Therefore, modifying the ET condition (3.4):
@@
$B_{n}^{Sub}(\bm{\alpha},\bm{\beta})=\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{K}\left(\frac{A_{i}}{n_{k1}/n_{+}}-\frac{1-A_{i}}{1-n_{k1}/n_{+}}\right)\left(Y_{i}-m_{A_{i}}(\bm{X}_{i};\bm{\beta})\right){\rm I}_{k}(\bm{X}_{i}).$ (F.2)
In fact, using (F.2), the proposed Bayesian procedure also achieves (approximately) double robustness, based on the same discussion presented in the main manuscript.